### v1

# Memory in the Burst Occurrence of Repeating FRBs

Abstract
Understanding the nature of repeating FRBs is crucial to probe the physics of FRBs. In this work, we analyze the statistics of waiting time between bursts of three repeating FRBs from four data sets. We find a universal pronounced dependency of the waiting times on the previous time interval (denoted as $\lambda_{0}$). We observe a temporal clustering where short waiting times tend to be followed by short ones, and long by long. This memory dependency is manifested in the conditional mean waiting time as well as in the conditional mean residual time to the next burst, both of which increase in direct proportion to $\lambda_{0}$. Consequently, the likelihood of experiencing a subsequent FRB burst within a given time window after the preceding burst is significantly influenced by the burst history. We reveal that, for the first time, these memory effects are present in the scale-invariant preconditioned waiting time distribution. We show that the memory effect provides a unified description of waiting times which may account for both the repeating FRBs and the apparent non-repeating FRBs (i.e. only observed one time). These results shed new light on the mechanism of FRBs.
Radio bursts (1339) — Radio transient sources (2008)
1 Introduction
Fast Radio Bursts (FRBs) represent one of the most powerful astronomical phenomena discovered in recent years (Cordes & Chatterjee, 2019; Bailes, 2022). Characterized by their brief yet energetic emissions, lasting merely a few milliseconds, they release energy comparable to what the sun emits over several days. The study of FRBs holds the promise of unveiling critical insights into the cosmic medium, the distribution of matter among galaxies and galaxy clusters (Petroff et al., 2019; Akahori et al., 2016). Moreover, observing the delay and frequency drift in FRBs offers valuable insights into the nature and origin of these mysterious cosmic phenomena (Hessels et al., 2019; Metzger et al., 2019), which can help constrain theoretical models and shed light on the mechanisms responsible for their generation. The research in FRBs also propels us into the realm of exploring unknown physical phenomena, including supernova explosions, neutron star mergers, black hole activities (Cordes & Chatterjee, 2019), and even exotic possibilities like axion miniclusters (Rosa & Kephart, 2018). These investigations bear significant physical implications and contribute to our understanding of fundamental astrophysical processes. One of the most significant discoveries in the study of FRBs is the identification of a repeating FRB known as FRB 121102 (Spitler et al., 2016). This repeating burst has provided valuable insights into the physical nature of FRBs and has supported models that propose an extragalactic neutron star as the origin of these bursts.
The origin and propagation mechanisms surrounding FRBs, however, remain mysterious, inspiring a abundance of theories. Among the proposed models, the collision or merger of compact objects like neutron stars or black holes, and the activity of highly magnetized neutron stars, known as magnetars, have attracted substantial attention. The discovery of X-ray emissions accompanying an FRB from a magnetar in the Milky Way, for instance, hints at the plausible role of magnetars in FRB production (Bochenek et al., 2020; Zhang, 2020; Li et al., 2021a). Furthermore, the young, highly magnetized, extragalactic neutron star model, substantiated by the observation of a repeating FRB known as FRB 121102, has also found support within possible models (Spitler et al., 2016).
Despite the progress made in understanding their physical properties through the study of host galaxies, the existing models do not fully explain all aspects of FRBs (Zhang, 2020, 2023). Particularly, the phenomenon of repeating bursts - a scenario where multiple bursts emerge from the same source - poses a substantial enigma. The speculation around the nature of repeating bursts suggests a contrast; while some FRBs exhibit repeating behavior, others appear as one-off events, with no subsequent activity detected. The presence of non-repeating FRBs cannot be entirely ruled out, yet the majority of observed non-repetitive bursts may likely still be repetitive bursts (Zhang, 2023). The current unseen apparent non-repeating reasons, along with the intermittent phenomena and extended time intervals between these FRBs as exhibited in existing records, should require further investigation.
Since the identification of the first FRB in 2007 (Lorimer et al., 2007), the tally of discovered FRBs has increased to over 700, with more than 60 shown repeating burst phenomena (Xu et al., 2023). The ongoing debate on the fundamental difference, if any, between the origins and physical mechanisms of repeating and non-repeating bursts continues to stimulate investigation. Recent efforts by FAST, involving continuous observations of several repeating FRBs with low threshold and high time resolution, have collected a wealth of data on repeating FRBs. This rich repository of information provides an opportunity to uncover more clues about the origin mechanisms of FRBs and the key temporal characteristics of repeating bursts through rigorous statistical analysis.
The temporal analysis of repeating FRBs suggests an evolving nature of their sources over time. If the emission of FRBs followed a Poisson process, the occurrence of bursts would be random, with a consistent number within any given time frame. Nevertheless, recent research has considered the likelihood of periodicity in the timing of FRB events (The CHIME/FRB Collaboration et al., 2022; Wei et al., 2022), and evidence points to a non-Poisson distribution in some repeating FRBs (Oppermann et al., 2018; Wang et al., 2023; Du et al., 2023). A newly identified characteristic, demonstrated by the scaling properties in burst timing across diverse sources, persists despite the variable environments of these sources (Du et al., 2023). The uniformity observed in time interval distributions from different FRB sources, even when influenced by varying fluences, emphasizes the significance of inter-event time statistics in understanding the dynamics of bursts. Contrasting with the distribution of burst energy, these findings could underline the potential to uncover key dynamics of the burst mechanism through statistical examination of the event time distributions (Kumar et al., 2020).
The specific distribution function, scaled by the average interval time of bursts $\bar{\lambda}$, is expressed as
$P_{I}(\lambda)\sim\frac{1}{\bar{\lambda}_{I}}f(\lambda/\bar{\lambda}_{I}).$ (1)
The mean waiting time, denoted as $\bar{\lambda}$ is dependent on magnitude $I$ of fluence and defined during continuous observing periods. This distribution closely approximates the function commonly known as the generalized Pareto distribution (Hosking & Wallis, 1987),
$f(\tau)\sim(1+\frac{\tau}{\beta})^{-\gamma},$ (2)
where $\beta$ and $\gamma$ are fitting parameters.
Based on the previously established scale-invariant waiting time distribution Eq. (1) (Du et al., 2023), this study explores further into the temporal characteristics of FRBs occurrences contained in $P$ (or $f$). Specifically, we examine the distribution of time intervals between successive bursts in repeating FRBs when the intervals between the last two bursts are $\lambda_{0}$, effectively the conditional probability distribution $P_{I}(\lambda|\lambda_{0})$. For Poisson processes, which lack memory effects, the outcome of $P_{I}(\lambda|\lambda_{0})$ is identical to $P_{I}(\lambda)$ and is independent of the value of $\lambda_{0}$. Our analysis reveals that $P_{I}(\lambda|\lambda_{0})$ is strongly dependent on the $\lambda_{0}$; smaller time intervals are more likely to be followed by smaller time intervals, and vice versa for larger time intervals. This clustering behavior in repeating FRBs is further quantified in various aspects throughout this paper, such as the average time intervals and residual time under the condition $\lambda_{0}$. Moreover, if one considers a scenario where the intervals between two consecutive bursts are $\lambda_{0}$, the results indicate that the memory effects in repeating FRBs significantly influence the probability of the next burst occurring within time $t$. This is crucial for assessing the event time of future bursts based on existing FRB observations.
2 Results
For the purpose of studying the issue of temporal clustering of repeating FRBs, we search four high-quality datasets, The first dataset is FRB 121102, comprising $1652$ bursts detected over $59.5$ hours across $47$ days (Li et al., 2021b). The second dataset focuses on FRB 20201124A, featuring $1863$ bursts observed over $82$ hours spanning $54$ days (Xu et al., 2022). The third dataset is related to FRB 20220912A, which include $1076$ bursts in $8.67$ hours over approximately $55$ days(Zhang et al., 2023). The fourth dataset involves another extremely active episode of FRB 20201124A, containing $881$ bursts observed in roughly $4$ hours within a $4$-day period(Zhou et al., 2022). For the purpose of differentiation, in this paper, the two datasets related to FRB 20201124A in sequence of time will be denoted as FRB 20201124A(A) and FRB 20201124A(B) in the following sections. All datasets were acquired by FAST recently and the high resolution detection has effectively reduced the flux threshold by up to three times compared to previous observations. Recent research on FRB 121102 and FRB 20201124A has revealed an absence of periodicity or quasi-periodicity on timescales ranging from $1$ ms to $1,000$ s (Li et al., 2021b; Xu et al., 2022). In this analysis, we investigate both waiting time and its scale invariance across different fluence thresholds. We uncover novel scaling and universality phenomena previously unreported in the field.
Due to the discontinuous nature of observations conducted by FAST, we redefine the waiting time as the recurrence interval between consecutive events observed within the same observational period, avoiding long gaps of observation [see Fig. 1(a)]. In our analysis, we concentrate on waiting times, denoted as $\lambda$, for bursts with fluence exceeding a specified threshold value $I$. Here, $\lambda=t_{i+1}-t_{i}$, where $t_{i}$ and $t_{i+1}$ represent the arrival times of the $i$th and $(i+1)$th burst, respectively.
As a typical example, the sequence of waiting times for FRB 121102 is shown in Fig. 1(b), where the horizontal line is the median waiting time. Pronounced clusters of both short and long waiting times, relative to the median, are evident. In contrast, Fig. 1(c) displays the sequence of waiting times after shuffling the original sequence, and it notably lacks the similar clustering behavior observed in Fig. 1(b). The clusters in Fig. 1(b) qualitatively illustrate the presence of a memory effect in the waiting time sequence, where longer waiting times typically follow longer ones, and shorter times follow shorter ones.
Figure 1: (a) The FAST telescope observed the sequence of FRB 121102 with a fluence greater than 0.2 Jy ms during continuous observation sessions.(b) Sequence of the waiting time above and below the median $\lambda$ = 120 s. (c) Same as (b) but for shuffled sequence of original data.
To quantitatively characterize the memory effect, we examine the distribution of subsequent waiting times under different preceding conditions, i.e., different values of the previous waiting time denoted as $\lambda_{0}$, represented as $P_{I}(\lambda|\lambda_{0})$. To distinguish the magnitude of $\lambda_{0}$, the entire waiting time sequence is sorted in ascending order and divided into four subsets $q_{1},q_{2},q_{3}$ and $q_{4}$, each containing an equal number of elements. Fig. 2 provides a comparative analysis of the waiting time distribution $\bar{\lambda}P_{I}(\lambda|\lambda_{0})$ when the preceding condition $\lambda_{0}$ is placed in $q_{1}$ (right green triangle) and $q_{4}$ (left blue triangle). Additionally, the distribution of waiting times without considering the preceding condition, $\bar{\lambda}P(\lambda)$, is plotted with the complete data sequence designated as $q_{0}$ (red dots). Moreover, different fluence thresholds were selected for different FRB sources, and the actual results are independent of the magnitude of the fluence threshold. It is worth noting that a secondary peak in the waiting time, ranging from a few ms to tens of ms, was observed in all four burst source datasets, potentially due to the specific criteria used to characterize separate outbursts.(Li et al., 2021b; Jahns et al., 2023; Xu et al., 2022; Zhang et al., 2023; Zhou et al., 2022). Unlike the first peak, the location of the secondary peak remains constant as the fluences increase, disrupting the universal scaling relation (Li et al., 2021b). For the occurrence of the first peak is depend on the intensity of activity from the originating source and to ensure a uniform treatment across all distributions, we only consider the waiting times within the scaling region; that is, the minimal waiting time interval surpasses the secondary peak range, approximately $10$s for FRB 121102 and 20201124A(A), and $1$s for FRB 20220912A and 20201124A(B).
Figure 2: The preconditioned distribution $P(\lambda|\lambda_{0})$ characterizes the waiting times $\lambda$ between successive FRBs that exceed specified thresholds for various sources. The preceding condition values $\lambda_{0}$ are taken from both the first (minimal) and last (maximum) quartiles, denoted as $q_{1}$ and $q_{4}$, respectively. For the sake of uniform notation, $q_{0}$ is used to represent the entire dataset, implying that $P(\lambda|\lambda_{0}\in q_{0})=P(\lambda)$. (a) $P(\lambda|\lambda_{0})$ for FRB 121102 with fluence thresholds of 0, 0.028 and 0.1 Jy ms; (b) $P(\lambda|\lambda_{0})$ for FRB 20201124A(A) with fluence thresolds of 0, 0.5, 1.0 Jy s; (c) $P(\lambda|\lambda_{0})$ for FRB 20220912A with fluence thresholds of 0, 0.1, 0.6 Jy ms; (d) $P(\lambda|\lambda_{0})$ for FRB 20201124A(B) with fluence thresholds of 0, 0.1, 0.5 Jy ms.
As illustrated in Fig. 2, $P_{I}(\lambda|\lambda_{0})$ exhibits a strong dependence on the preceding condition values $\lambda_{0}$, and changes its functional form accordingly. Similar to $P_{I}(\lambda)$, $P_{I}(\lambda|\lambda_{0})$ demonstrates independence from the magnitude of fluence $I$. Hence, in subsequent analyses, the magnitude of fluence $I$ will be omitted unless explicitly needed. As observed from Fig. 2, when $\lambda_{0}$ is taken from $q_{1}$ and $q_{4}$ respectively, shorter waiting times are more probable ($P(\lambda|\lambda_{0})$) in the interval $\lambda/\bar{\lambda}<1$ compared to longer waiting times. Conversely, for $\lambda/\bar{\lambda}>1$, the situation is reversed. Across the entire range, the $P(\lambda|\lambda_{0})$ without any preceding condition ($\lambda_{0}\in q_{0}$) falls between theses two cases. Further analysis reveals that $P(\lambda|\lambda_{0})$ can also be well approximated by $P(\lambda|\lambda_{0})\sim f(\lambda/\bar{\lambda},\lambda_{0}/\bar{\lambda})/\bar{\lambda}$. Deviations from an exponential function further demonstrates the strong memory effects that the preceding condition values $\lambda_{0}$ has on subsequent FRB bursts and their degree of clustering. The parameter $\gamma$ in Eq. (2) describes these effects, with more details provided in Table 1.
Table 1: Values of the parameter $\gamma$ of the waiting time distribution $f(\lambda|\lambda_{0})$, preconditioned or not. $q_{1}$ and $q_{4}$ means $\lambda_{0}$ is from the first and the last part of sorted data, while $q_{0}$ is the unconditional case.
FRB 121102 FRB 20220912A FRB 20201124A(A) FRB 20201124A(B) Average
$q_{1}$ 2.40 3.07 2.55 3.88 $2.98\pm 0.65$
$q_{4}$ 2.08 2.06 2.13 2.07 $2.09\pm 0.41$
$q_{0}$ 2.22 2.63 2.40 3.04 $2.57\pm 0.38$
To quantify the impact of different preceding condition $\lambda_{0}$ on subsequent burst events, we proceed to calculate the average value of the following burst waiting time distribution, denoted as $\hat{\lambda}(\lambda_{0})$. From Eq. (1), we obtain a fluence independent functional form $\hat{\lambda}(\lambda_{0})\sim\bar{\lambda}g(\lambda_{0}/\bar{\lambda})$. Fig. 3 shows the ratio $\hat{\lambda}(\lambda_{0})/\bar{\lambda}$ as a function of $\lambda_{0}/\bar{\lambda}$. Due to the strong memory effect in the system, the value of $\hat{\lambda}/\bar{\lambda}$ is significantly below one when $\hat{\lambda}/\bar{\lambda}$ is much below one, and notably above one when $\hat{\lambda}/\bar{\lambda}$ is much above one. This implies that shorter waiting times are more likely to be followed by shorter ones, and longer waiting times by longer ones. For comparison, when the same preceding condition $\lambda_{0}$ is applied after random shuffling, the value of $\hat{\lambda}/\bar{\lambda}$ is approximately 1, indicating that the memory effect has been disrupted.
Figure 3: The average value of conditional probability of waiting time $\hat{\lambda}(\lambda_{0})/\bar{\lambda}$. They have similar form as $\lambda_{0}/\bar{\lambda}$ increasing with the same fluence thresholds shown in Fig. 2. The symbols of plus correspond to the shuffled sequences of waiting times which are fluctuating across the $\hat{\lambda}(\lambda_{0})/\bar{\lambda}\approx 1$, indicative of a lack of memory in the data.
The obtained $\hat{\lambda}(\lambda_{0})$ reflects the average time to wait for the next FRB burst under different preceding condition values $\lambda_{0}$. If we want to know more detailed temporal information about the next burst when the time interval between the previous two bursts is $\lambda_{0}$, this can be obtained by calculating the probability of another burst occurring within a certain time (or risk function), denoted as $R(t|\lambda_{0})$. Fig. 4 compares the probability distribution of the timing of the next burst when the time interval $\lambda_{0}$ between the previous two bursts is either minimal ($\in q_{1}$) or maxmal ($\in q_{4}$). For FRB 121102 (Fig. 4(a)) and FRB 20220912A (Fig. 4(b)), the value of $\lambda_{0}$ has a significant impact on the timing of the next burst within a certain time range. As a comparison, for the same burst source FRB 20201124A(A) and FRB 20201124A(B) during different active periods, $R(t|\lambda_{0})$ is relatively less affected by the magnitude of the preceding condition values $\lambda_{0}$, but maintains the property that a smaller $\lambda_{0}$ is followed by a shorter time until the next FRB burst, and a larger $\lambda_{0}$ is followed by a relatively longer waiting time for the next burst. These characteristics are important in understanding the bursting nature of FRBs.
Figure 4: The conditional probability of waiting time $R(t|\lambda_{0})$, exceeding a specified fluence threshold and separated by time $\lambda_{0}$, are succeeded by the next FRB within time t. The $\lambda_{0}$ comes from either $q_{1}$ or $q_{4}$ and thresholds are 0.1, 0.6 and 0.5 Jy ms for FRB 121102, FRB 20220912A and FRB 20201124A(B) and 1.0 Jy s for FRB 20201124A(A).
From the preceding analysis, it is evident that the time intervals between the first two bursts in repeating FRBs exert a significant influence on the occurrence of subsequent bursts. A realistic question arises: if a time interval $t$ has elapsed since the last burst without a subsequent occurrence, what is the average waiting time until the next burst? To answer this, one must calculate the expected residual time $\hat{\lambda}(t|\lambda_{0})$. For the special case when $t=0$, $\hat{\lambda}(t=0|\lambda_{0})=\hat{\lambda}(\lambda_{0})$. Based on the definition of residual time and the known waiting time distribution (Eq. 2), one can deduce that
[…]
For distribution $P(\lambda|\lambda_{0})$ that satisfied the Eq. (2), we can obtain
$\hat{\lambda}(t|\lambda_{0})\sim\frac{(t+\beta)(\gamma-1)}{\gamma^{2}-3\gamma+2},$ (4)
where the integral is divergence when $\gamma$ is equal to $1.0$ and $2.0$.
When there is no memory effect, $\hat{\lambda}(t|\lambda_{0})$ is independent of both $\lambda_{0}$ and $t$. Conversely, if a memory effect exists, a clear dependency will be observed. To quantitatively characterize the dependency of $\hat{\lambda}$ on $\lambda_{0}$ and $t$, the following scenarios were considered: firstly, intervals $q_{+}$ and $q_{-}$ that are respectively greater and smaller than its median, and $\lambda_{0}$ is taken from intervals $q_{+}$ and $q_{-}$, denoted as $\lambda^{+}_{0}$ and $\lambda^{-}_{0}$. Additionally, $\hat{\lambda}(t)$ without any preceding condition is also considered; secondly, For comparison, $t$ is set to either 00 or $\bar{\lambda}/2$. The results are shown in Fig. 5, where sample A, B, C, D represent the datasets for FRB 121102, FRB 20220912A, FRB 20201124A(A), and FRB 20201124A(B), respectively. For the case of $t=0$ (Fig. 5(a)), since $\hat{\lambda}(0)=(\hat{\lambda}(0|\lambda^{-}_{0})+\hat{\lambda}(0|\lambda^{+}_{0}))/2$, the results show that $\hat{\lambda}(0|\lambda^{-}_{0})$ (Blue square) and $\hat{\lambda}(0|\lambda^{+}_{0})$ (Blue left triangle) are approximately symmetrical about $\hat{\lambda}(0)$ (Green solid circle) $=1$, and due to the memory effect, $\hat{\lambda}(0|\lambda^{+}_{0})$ is significantly greater than $\hat{\lambda}(0|\lambda^{-}_{0})$. Similar memory effects as shown in Fig.5(a) are also evident in Fig. 5(b), i.e., $\hat{\lambda}(\frac{\bar{\lambda}}{2}|\lambda^{+}_{0})$ is significantly greater than $\hat{\lambda}(\frac{\bar{\lambda}}{2}|\lambda^{-}_{0})$. Furthermore, due to the memory effect, as the time elapsed is $t=\lambda/2$ since the last burst was introduced, the symmetry between the residual times as shown in Fig. 5(a) is broken due to the arising of $\bar{\lambda}$, which is mostly contributed from $q_{+}$.
Figure 5: The conditional residual time $\hat{\lambda}(t|\lambda_{0})$. The $\lambda_{0}$ is either from $q_{+}$ (above median of entire data) or $q_{-}$ (below median of entire data). Two cases are considered for (a) $t=0$ and (b) $t=\bar{\lambda}/2$. Datasets A, B, C, and D correspond to FRB 121102, FRB 20220912A, FRB 20201124A(A), and FRB 20201124A(B), respectively. Fluences thresholds are the same as in Fig. 4.
As illustrated in Fig. 5, after a time interval $t$ has elapsed since the last burst, the expected residual time to the next burst is generally greater than when $t=0$. This can be directly inferred from Eq. 4. Generally, as $t$ increases, $\hat{\lambda}(t|\lambda_{0})$ also monotonically increases, with the rate of increase influenced by the value of $\gamma$. This behavior contrasts with the memoryless nature of the Poisson process (Sornette & Knopoff, 1997; Corral, 2005; Livina et al., 2005).
3 Discussion
Contrary to intuitive view, the probability of next burst occurring does not increase with the time $t$ elapsed since the last burst; rather, it decreases and the expected residual time to the next burst $\hat{\lambda}$ is also increased. This serves as a more direct characterization of temporal clustering, where waiting times exhibit a tendency to cluster together.
An intuitive explanation for this phenomenon shown that when a significant amount of time has elapsed since the last FRB burst, the system enters a enhanced ’period of inactivity,’ during which the likelihood of an extended waiting time increases. It is important to note, however, that these periods of inactivity are not fundamentally distinct from other waiting times, as they are all governed by the same continuous distribution. Furthermore, the fluence of the subsequent burst does not increase with the waiting time but appears to be independent of it.
Recent research, as referenced in (Aschwanden & McTiernan, 2010), offers a generalized understanding of nonstationary Poisson processes, which maybe helpful in interpreting the results. Specifically, when the variability of the burst rate follows a gradual pattern and can be described by piecewise exponential functions, the exponent $\gamma$ in the waiting time distribution approximates 3.0. Conversely, if the burst rate variability exhibits sharp, intermittent spikes like to $\delta$-functions, the exponent $\gamma$ in the waiting time distribution tends toward 2.0. This behavior, commonly referred to as ”clusterization”, finds also in earthquake and solar flare statistics (Corral, 2004, 2005; Livina et al., 2005; Baiesi et al., 2006; Aschwanden & McTiernan, 2010), which serves as indirect evidence for non-trivial space-time correlations among events. Consequently, the value of the exponent $\gamma$ serves as a critical indicator for determining whether the burst rate undergoes gradual or intermittent fluctuations. As a result, the bursts themselves are not periodic, the activity may occur within periodic windows (Wei et al., 2022). Some recent studies have sought to distinguish between periodic windowed behavior and mere event clusters through time clustering analysis (Denissenya et al., 2021).
In this study, the analysis of four datasets from three FRB sources reveals that the value of $\gamma$ in the waiting time distribution (Eq. (2)) ranges approximately between $2.2$ and $3.0$ (Table 1). For cases where $\gamma$ is close to $2.0$, we made an estimate based on Eq. (4), when $\beta=1.0,\gamma=2.1$, the expected residual time to the next burst $\hat{\lambda}$ is roughly 10 times the magnitude of the elapsed time $t$ after scaling. Here, the actual value of each scale $t$ is related to the average waiting time under different fluence thresholds. If $\beta$ remains constant and $\gamma$ decreases to $2.01$, then $\hat{\lambda}$ will be roughly $100$ times $t$. As $\gamma$ gradually approaches $2.0$, $\hat{\lambda}$ will become larger. The three burst sources analyzed in this paper have very different burst environments and properties. However, the statistical analysis results of these three repeating FRBs with the largest sample sizes indicate that the scale invariant waiting time distribution and the clustering and memory of bursts are universal. This may be well used to explain the burst event characteristics of repeating FRBs that have been active in the past, as well as to explain the burst event characteristics that have only been observed once but may be repeating bursts. Although the existence of true one-offs is not ruled out, Eq. (4) has also shown that when $\gamma=2.0$, the expected residual time diverges and tends to be infinite, and this corresponds to the case of spikes like $\delta$ function of time dependent burst rate (Aschwanden & McTiernan, 2010), which mimics a true one-off burst from an observational perspective. Therefore Eq. (2) actually provides a unified description of the statistical behavior of the waiting time distribution, satisfying this equation for both repeating FRBs and one-off FRBs.
The complexity of waiting time between FRBs suggests that the physical models proposed thus far may not fully capture the complexity of the phenomenon. Currently, over 700 FRBs have been observed, of which more than 60 exhibit repeating bursts (Xu et al., 2023). Some analyses suggest that one-off and repeating fast radio bursts have some statistical differences (Chen et al., 2022; Cui et al., 2020). There are also studies discussing whether repeating and non-repeating FRBs have the same burst mechanisms (Caleb et al., 2019; Zhang, 2020, 2023), and partially explaining the statistical differences between one-off and repeating FRBs from an observational perspective (Iwazaki, 2021). However, more research is needed to determine the precise mechanisms that lead to the intermittent nature of FRBs and the extended periods between bursts. The ubiquity of this behavior implies the presence of a straightforward mechanism. In this mechanism, as time progresses, the variable responsible for triggering the burst increasingly deviates from the burst threshold, on average. These ’fluctuations’ in the variable effectively preserve the memory of the last event within the system for extended periods, thereby contributing to the observed absence of time dependent aging. It is important to emphasize that these findings are in fact model independent, meaning they do not rely on any preconceived model of burst occurrence. Bak et al. (Bak, 1996; Bak et al., 2004) suggests that the observed scaling law in burst events reflects a complex spatiotemporal pattern. This may be interpreted as intermittent energy release from a self-organized system , which is consistent with the theoretical constructs of critical phenomena. Analogous memory effects, attributable to long-term correlations, have also been recently identified in the statistical analysis of two FRBs (Wang et al., 2023), which address the self-similarity inherent in these processes, the framework of SOC currently offers the most compelling approach.
4 Summary
In conclusion, we have analyzed the statistics of waiting times between bursts (above certain fluence threshold values $I$) of three repeating FRBs. Our observations reveal a notable memory effect in these events, wherein small waiting times tend to follow small ones, and large waiting times follow large ones. This clustering of FRBs has been quantified using four metrics: (i) the conditional distribution function, (ii) the conditional mean recurrence time, (iii) the conditional risk function, and (iv) the conditional residual time. Each of these metrics significantly deviates from their respective unconditional counterparts. The memory effect proves particularly beneficial as it enables us, as demonstrated in Fig. 5, to utilize relevant elapsed time data since last burst to achieve a more accurate burst estimation. For both repeating FRBs in their active periods and the apparent non-repeating FRBs that have been observed to burst only once, a unified description of waiting time is provided.
Data availability
All data analysed in this paper are included in references (Li et al., 2021b; Xu et al., 2022; Zhang et al., 2023; Zhou et al., 2022) separately. They are also openly available in Science Data Bank at https://doi.org/10.11922/sciencedb.01092 (Li et al., 2021b), https://cstr.cn/31253.11.sciencedb.08058 (Zhang et al., 2023), https://cstr.cn/31253.11.sciencedb.06762 (Zhou et al., 2022), and at the PSRPKU website, https://psr.pku.edu.cn/index.php/publications/frb20201124a/ (Xu et al., 2022).
Acknowledgements: We would like to thank Mingyu Ge and Sai Wang for useful discussions. The authors acknowledge the computing resources provided by National High Energy Physics Data Center and High Energy Physics Data Center, Chinese Academy of Science. S.-L.X. acknowledge the support by the National Natural Science Foundation of China (Grant No. 12273042), S.X. acknowledge the support by the National Natural Science Foundation of China (NO. 12303043), and C.C. acknowledge the support by the Science Foundation of Hebei Normal University (No. L2023B11), the National Natural Science Foundation of China (No. 12303045) and Natural Science Foundation of Hebei Province (No. A2023205020).